### v1

@@ Abstract
In soccer video analysis, player detection is essential for identifying key events and reconstructing tactical positions. The presence of numerous players and frequent occlusions, combined with copyright restrictions, severely restricts the availability of datasets, leaving limited options such as SoccerNet-Tracking and SportsMOT. These datasets suffer from a lack of diversity, which hinders algorithms from adapting effectively to varied soccer video contexts. To address these challenges, we developed SoccerSynth-Detection, the first synthetic dataset designed for the detection of synthetic soccer players. It includes a broad range of random lighting and textures, as well as simulated camera motion blur. We validated its efficacy using the object detection model (Yolov8n) against real-world datasets (SoccerNet-Tracking and SportsMoT). In transfer tests, it matched the performance of real datasets and significantly outperformed them in images with motion blur; in pre-training tests, it demonstrated its efficacy as a pre-training dataset, significantly enhancing the algorithm’s overall performance. Our work demonstrates the potential of synthetic datasets to replace real datasets for algorithm training in the field of soccer video analysis.
@@ 2.3 Validation
To validate the effectiveness of our method, we generated 37,621 normal and 17,052 motion-blurred images using the simulator, dividing them into training and validation datasets at a 4:1 ratio, followed by conducting transfer learning and pretraining experiments. Both of the experiments were conducted with an image size of 640x640, a learning rate of 0.01, for up to 300 epochs, and an Intersection over Union (IoU) threshold of 0.7, with data augmentation applied. To ensure consistency with the actual annotations, we manually expanded the horizontal width of the detection boxes by two pixels. In both experiments, we utilized the YOLOv8n model [21] with the pretrained parameters “yolov8n.pt”, noted for its minimal parameter count. In both experiments, we evaluated the performance of the YOLOv8n model using Average Precision (AP), a metric that balances precision (the model’s ability to correctly identify relevant objects) and recall (the model’s ability to detect all relevant objects). A higher AP score reflects greater model accuracy. Specifically, AP50 measures AP at an IoU threshold of 0.5, providing insight into the model’s basic detection capabilities. In contrast, AP50:5:95 calculates the average AP across multiple IoU thresholds (ranging from 0.5 to 0.95 in increments of 0.05), offering a comprehensive evaluation of the model’s performance under varying detection strictness. In the fields of autonomous driving and pedestrian tracking, models trained on synthtic datasets have demonstrated superior generalization capabilities compared to those trained on real datasets [31, 28]. Therefore, we decided to conduct transfer experiments to evaluate the generalization performance of the model trained on SoccerSynth-Detection.
@@ 3 Results
First, in the transfer experiment, we evaluated the generalization performance of the model trained on SoccerSynth-Detection. TAs illustrated in Table 1, while the Yolov8n model trained on the SoccerSynth-Detection dataset exhibited a minor decline of 2.2 percentage points at AP50, it demonstrated a gain of 1.0 percentage points at AP50:5:95. Furthermore, performance evaluations on the SoccerNet-Tracking dataset indicated substantially better outcomes. The Yolov8n model, when trained on the SoccerSynth-Detection dataset, achieved significant enhancements of 6.6 percentage points at AP50 and 4.5 percentage points at AP50:5:95.
@@
Next, in the pretrainig experiment, we examined the performance in settings with limited data. TTable 2 demonstrates that when using SoccerNet-Tracking as the real-world dataset, the YOLOv8n model pretrained with SoccerSynth-Detection achieved an improvement of 0.1 to 1.5 percentage points in AP50 and AP50:5:95 metrics, effectively mitigating the adverse impact caused by reduced data volume. When SportsMOT was used as the real-world dataset with a subsampling rate of 1/8, the pretrained YOLOv8n model achieved the highest AP50 of 95.3 percent.

### v2

@@ Abstract
In soccer video analysis, player detection is essential for identifying key events and reconstructing tactical positions. The presence of numerous players and frequent occlusions, combined with copyright restrictions, severely restricts the availability of datasets, leaving limited options such as SoccerNet-Tracking and SportsMOT. These datasets suffer from a lack of diversity, which hinders algorithms from adapting effectively to varied soccer video contexts. To address these challenges, we developed SoccerSynth-Detection, the first synthetic dataset designed for the detection of synthetic soccer players. It includes a broad range of random lighting and textures, as well as simulated camera motion blur. We validated its efficacy using the object detection model (Yolov8n) against real-world datasets (SoccerNet-Tracking and SportsMoT). In transfer tests, it matched the performance of real datasets and significantly outperformed them in images with motion blur; in pre-training tests, it demonstrated its efficacy as a pre-training dataset, significantly enhancing the algorithm’s overall performance. Our work demonstrates the potential of synthetic datasets to replace real datasets for algorithm training in the field of soccer video analysis. The dataset is available at https://github.com/open-starlab/SoccerSynth-Detection.
@@ 2.3 Validation
To validate the effectiveness of our method, we generated 37,621 normal and 17,052 motion-blurred images using the simulator, dividing them into training and validation datasets at a 4:1 ratio, followed by conducting transfer learning and pretraining experiments. Both of the experiments were conducted with an image size of 640x640, a learning rate of 0.01, for up to 300 epochs, and an Intersection over Union (IoU) threshold of 0.7, with data augmentation applied. To ensure consistency with the actual annotations, we manually expanded the horizontal width of the detection boxes by two pixels. In both experiments, we utilized the YOLOv8n model [21] with the pretrained parameters “yolov8n.pt”, noted for its minimal parameter count. In both experiments, we evaluated the performance of the YOLOv8n model using Average Precision (AP), a metric that balances precision (the model’s ability to correctly identify relevant objects) and recall (the model’s ability to detect all relevant objects). A higher AP score reflects greater model accuracy. Specifically, AP50 measures AP at an IoU threshold of 0.5, providing insight into the model’s basic detection capabilities. In contrast, AP50:5:95 calculates the average AP across multiple IoU thresholds (ranging from 0.5 to 0.95 in increments of 0.05), offering a comprehensive evaluation of the model’s performance under varying detection strictness.
In the fields of autonomous driving and pedestrian tracking, models trained on synthtic datasets have demonstrated superior generalization capabilities compared to those trained on real datasets [31, 28]. Therefore, we decided to conduct transfer experiments to evaluate the generalization performance of the model trained on SoccerSynth-Detection.
@@ 3 Results
First, in the transfer experiment, we evaluated the generalization performance of the model trained on SoccerSynth-Detection. As illustrated in Table 1, while the Yolov8n model trained on the SoccerSynth-Detection dataset exhibited a minor decline of 2.2 percentage points at AP50, it demonstrated a gain of 1.0 percentage points at AP50:5:95. Furthermore, performance evaluations on the SoccerNet-Tracking dataset indicated substantially better outcomes. The Yolov8n model, when trained on the SoccerSynth-Detection dataset, achieved significant enhancements of 6.6 percentage points at AP50 and 4.5 percentage points at AP50:5:95.
@@
Next, in the pretrainig experiment, we examined the performance in settings with limited data. Table 2 demonstrates that when using SoccerNet-Tracking as the real-world dataset, the YOLOv8n model pretrained with SoccerSynth-Detection achieved an improvement of 0.1 to 1.5 percentage points in AP50 and AP50:5:95 metrics, effectively mitigating the adverse impact caused by reduced data volume. When SportsMOT was used as the real-world dataset with a subsampling rate of 1/8, the pretrained YOLOv8n model achieved the highest AP50 of 95.3 percent.
@@ 5 Conclusion
Author Contribution H.Q. implement the system and H.Q. and K.F. wrote the main manuscript text. All authors discussed this research and reviewed the manuscript.